OpenStudy (anonymous):

The sequence is recursive. Find the value of the next term in the sequence. 15, 26, 48, 92, . .

OpenStudy (helder_edwin):

I think it is \[\large a_n=a_{n-1}\cdot2-4 \] for n>1.

OpenStudy (anonymous):

26-15=11=11*1 48-26=22=11*2 92-48=44=11*4 92+11*8=?

OpenStudy (eyust707):

lol @surjithayer I didn't notice that

OpenStudy (eyust707):

The whole thing with recursive relationships is they depend on the previous term. I feel like what @helder_edwin is saying helps to understand them a little better. Basically what he's saying is the the next term is simply the previous term *2 -4

OpenStudy (eyust707):

15*2 -4 = 26 26 *2 - 4 = 48 48*2 - 4 = 92 92*2 - 4 =

OpenStudy (anonymous):

\[a _{1}=15\] \[a _{n+1}=15+11\left( 2 \right)^{n-1}\] \[n \ge 1\]or n is a natural number.